OpenStudy (anonymous):

Rickey has fraction 7 over 8 of a cup of cereal in a bowl. He then gave two fraction 1 over 4 cups of cereal from his bowl to his little sister. Which of the following best explains the amount of cereal left in Rickey's bowl? Please help

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just subtract it

OpenStudy (anonymous):

7/8 - 1/4

OpenStudy (anonymous):

ok. working it

OpenStudy (anonymous):

3/4?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

7/8 - 1/4 = 7/8 - 2/8= ??

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thanks @kliedako

OpenStudy (anonymous):

The graph below shows the fine that a DVD club member pays based on the number of days a DVD is overdue: I think that if the DVD is return it won't cost? any suggestion plz?